Laplace inverse using the convolution method

گربه به سیریل
Answer #1

xt { (54) Sett)} (S+1) CS2+4) IS L(F(S)) = f(t) , X EG()) = g(t) then It ( f(s)* Gls)) = 4( f (I) g(t-1)dt t (F(8) 46 (s)Lt ( f(s) & G (s)) = let [ etsin ał -2 et cas 2 t] set (sin 2tr. 2 cosat) +2 = sin 2t - 2 coset + et lo